• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 00:01:41 on April 08, 2008

    Entry number 230653

    keyword=Swing shift summary

    Swing shift summary.
    David Hamilton (SL)
    Nikos Sparveris (TO)
    Kevin Fissum (3rd)

    16:00 - shift started with no beam. MCC are tuning into Hall C.

    16:40 - MCC called to say that the magnet experts need another hour
    before beam can be restored to the hall.

    17:30 - Hall status changed to controlled access to allow Brad to go into
    the hall.

    18:15 - Brad finished installing a camera on the BB helium bag gas
    bubbler. Hall status is changed back to beam permit.

    18:20 - HRSl moved into position for elastic data-taking (35.5 degrees,
    969.3 MeV/c) for BB HV MWDC scan. Awaiting beam.

    19:00 - beam restored to hall.
    Target changed to 6cm LH2.
    BB is still at 390A and 54 degrees.
    Request 5 microamps with raster off.

    Run 2497 - 2499, 2501, 2502, 2510 were junk due to BB HV trip caused by
    bad beam alignment.
    Run 2500 - 1550V threshold +5V, 200k - replay ok
    Run 2503 - 1550V threshold +7.5V, 150k
    Run 2504 - 1550V threshold +10V, 150k

    Run 2505 - 1600V threshold +10V, 150k
    Run 2506 - 1600V threshold +7.5V, 150k
    Run 2507 - 1600V threshold +5V, 150k - replay ok

    Run 2508 - 1500V threshold +5V, 150k
    Run 2509 - 1500V threshold +7.5V, 150k
    Run 2511 - 1500V threshold +10V, 150k - replay ok

    Run 2512 - 1450V threshold +10V, 150k
    Run 2513 - 1450V threshold +7.5V, 150k
    Run 2514 - 1450V threshold +5V, 140k

    Run 2515 - 1400V threshold +5V, 150k - replay ok
    (comment line in header is wrong for 2515)
    Run 2516 - 1400V threshold +7.5V, 150k
    Run 2517 - 1400V threshold +10V, 150k

    21:30 - BB MWDC HV-threshold scan complete.
    BB field switched to 0A.
    0.5 microamps requested, raster off, fast FB off.
    Run 2518 - 1400V threshold +5V, 150k, BB I 0A
    Run 2520 - 1450V threshold +5V, 150k, BB I 0A

    22:00 - BB current switched back to 390A. Short BB MWDC HV/threshold scan
    with T3 events. 5 microamps.
    Run 2524 - 1550V threshold +5V, 670k, BB 390A, T3 - replay ok
    Run 2525 - 1550V threshold +7.5V, 500k, BB 390A, T3
    Run 2528 - 1550V threshold +10V, 590k, BB 390A, T3

    22.40 - Set beam current to 3 microamps, MWDC HV 1550V threshold 5V BB 390A.
    Run 2529 - 1550V threshold +5V, 0.29M, BB 390A, T3, current 3 microamps
    Run 2531 - 1550V threshold +5V, 0.27M, BB 390A, T3, beam 1 microamp

    23:00 - BB coil current set to 200A
    Run 2532 - 1550V threshold +5V, 0.3M, BB 200A, T3, beam 1 microamp

    23:10 - BB coil current set to 600A
    Run 2533 - 1550V threshold +5V, 0.3M, BB 600A, T3, beam 3 microamps

    23:20 - BB coil current set to 700A
    Run 2534 - 1550V threshold +5V, 0.3M, BB 700A, T3, beam 3 microamps

    23:50 - HRSl set to 1194 MeV/c (Q2 and Q3 cycled through 1400A) and 16.5
    degrees (kinematics 2).
    Target moved to Ta for elastic calibration data.
    Requested beam at 5 microamps, FFB on, raster off.

    Run 2535 - HRSl singles, Ta elastic run, 150k.

    Shift ended during move to C target for short elastic run.